# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 4.2.1 Intuitionistic non-deterministic valuations
Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is undetermined in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$. Hence, the failure of excluded middle (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$. Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly undetermined in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$. Hence, the failure of $\alpha\vee(\alpha\to\beta)$ (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val_{+}(\mathcal{M}_{IPL})$ (recall Definition 4.14) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
@@
Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$ for every $\alpha$. If $v_{\Delta}(\neg\alpha)=\mathbf{T}$ then $\neg\alpha\in\Delta$, by definition. Hence, $\alpha\notin\Delta$ and so $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Now, if $v_{\Delta}(\neg\alpha)=\mathbf{U}$ then $\alpha\notin\Delta$, by definition. Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Finally, if $v_{\Delta}(\neg\alpha)=\mathbf{F}$ then $\alpha\in\Delta$, by definition. Hence $v_{\Delta}(\alpha)=T$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
@@
– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
– If $\alpha\in\Delta$ then $\beta\in\Delta$, by MP. Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I2) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\notin\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I3) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{F}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\in\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)=\mathbf{T}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
(C1) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{T}$. Then $\alpha\in\Delta$ and $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(C2) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{F}$. Then, either $\alpha\notin\Delta$ or $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ or $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
(D1) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}$. Then, either $\alpha\in\Delta$ or $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=\mathbf{T}$ or $v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(D2) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{F}$. Then $\alpha\notin\Delta$ and $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a level valuation in $\mathcal{M}_{IPL}$ such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
@@
Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val_{+}(\mathcal{M}_{IPL})$, by Proposition 4.21.
[…]
(1) $\alpha=\neg\beta$ for some $\beta$. By Definition 4.19, $\neg\beta\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\neg\beta$. By the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\neg\beta$. By (Ax9) and by MP, $\Delta,\beta\to\beta,\beta\to\neg\beta\vdash_{IPL}\neg\beta$. But $\beta\to\beta\in\Delta$ and $\beta\to\neg\beta\in\Delta$, hence $\Delta\vdash_{IPL}\neg\beta$, that is, $\neg\beta\in\Delta$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\neg\beta$. Then, there exists a $\neg\beta$-saturated set $\Delta^{\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime}$. By IH, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime}}(\neg\beta)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime}$.
(2) $\alpha=\beta\to\gamma$ for some $\beta$ and $\gamma$. By Definition 4.19, $\beta\to\gamma\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\beta\to\gamma$. By MP, $\Delta,\beta\vdash_{IPL}\gamma$ and so, by the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\beta\to\gamma$. Then, there exists a $(\beta\to\gamma)$-saturated set $\Delta^{\prime\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime\prime}$. By IH, $v_{\Delta^{\prime\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime\prime}}(\beta\to\gamma)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime\prime}$.
@@
Theorem 4.23 (Completeness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
@@
Suppose that $\Gamma\nvdash_{IPL}\varphi$. Then, there exists a $\varphi$-saturated set $\Delta$ in IPL such that $\Gamma\subseteq\Delta$. By Proposition 4.22, $v_{\Delta}$ is a level valuation in $\mathcal{M}_{IPL}$ such that $v_{\Delta}(\alpha)=\mathbf{T}$ for every $\alpha\in\Gamma$, but $v_{\Delta}(\varphi)\neq\mathbf{T}$. This shows that $\Gamma\not\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
4.2.2 Intuitionistic partial (level) valuations and intuitionistic truth tables

### v4

@@ 8.2 Level valuations for $\mathcal{M}_{IPL}$
The observations about the meaning of $\mathbf{U}$ made after Definition 8.4 can be recast now. Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is unknown in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$. Hence, the failure of excluded middle $\alpha\vee\neg\alpha$ (a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$. Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly unknown in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$. Hence, the failure of Peirce’s law $\alpha\vee(\alpha\to\beta)$ (a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$. The function $v_{\Delta}$ can be seen as a 3-valued characteristic map associated to $\Delta$, in which the third value $\mathbf{U}$ reflects the undetermined status of a sentence that does not belong to $\Delta$. Only sentences of the form $\neg\alpha$ or $\alpha\to\beta$ can be unknown or undetermined (in this sense) w.r.t. a $\varphi$-saturated set in $\mathbf{IPL}$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val^{nd}_{+}(\mathcal{M}_{IPL})$ (recall Definition 8.5) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
@@
Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{IPL}(v_{\Delta}(\alpha))$ for every $\alpha$. If $v_{\Delta}(\neg\alpha)=\mathbf{T}$ then $\neg\alpha\in\Delta$, by definition. Hence, $\alpha\notin\Delta$ and so $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{IPL}(v_{\Delta}(\alpha))$. Now, if $v_{\Delta}(\neg\alpha)=\mathbf{U}$ then $\alpha\notin\Delta$, by definition. Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{IPL}(v_{\Delta}(\alpha))$. Finally, if $v_{\Delta}(\neg\alpha)=\mathbf{F}$ then $\alpha\in\Delta$, by definition. Hence $v_{\Delta}(\alpha)=\mathbf{T}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{IPL}(v_{\Delta}(\alpha))$.
@@
– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
– If $\alpha\in\Delta$ then $\beta\in\Delta$, by (DM). Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I2) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\notin\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 8.13. Then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I3) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{F}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\in\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 8.13. Then $v_{\Delta}(\alpha)=\mathbf{T}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
(C1) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{T}$. Then $\alpha\in\Delta$ and $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(C2) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{F}$. Then, either $\alpha\notin\Delta$ or $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ or $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
(D1) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}$. Then, either $\alpha\in\Delta$ or $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=\mathbf{T}$ or $v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(D2) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{F}$. Then $\alpha\notin\Delta$ and $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{IPL}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a level valuation in $\mathcal{M}_{IPL}$ such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
@@
Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val^{nd}_{+}(\mathcal{M}_{IPL})$, by Proposition 8.16.
Inductive step: Assume that, for every $\varphi^{\prime}$ and for every $\varphi^{\prime}$-saturated set $\Delta^{\prime}$, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{k}$ for every $k\leq n$, for a given $n\geq 0$ (Induction Hypothesis, IH). Let us prove that $v_{\Delta}\in\mathcal{L}^{IPL}_{n+1}$. By IH, $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$. Let $\alpha\in For(\Omega)$ such that $v_{\Delta}(\alpha)=\mathbf{U}$. By Definition 8.14, $\alpha\neq p$ for every propositional variable $p$; and $\alpha\neq\beta\#\gamma$ for $\in\{\land$$\vee\}$ and every formulas $\beta,\gamma$. Then, there are two cases to analyze:
(1) $\alpha=\neg\beta$ for some $\beta$. By Definition 8.14, $\neg\beta\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\neg\beta$. By the deduction metatheorem (DM), $\Delta\vdash_{IPL}\beta\to\neg\beta$, and so the sequent $\Delta_{0}\Rightarrow\beta\to\neg\beta$ is provable in LI, for some finite set $\Delta_{0}\subseteq\Delta$. But, for every finite multiset $\Gamma$, the sequent $\Gamma,\beta\to\neg\beta\Rightarrow\neg\beta$ is provable in LI. Indeed, from axiom $\Gamma,\beta\Rightarrow\beta$ it follows $\Gamma,\beta,\neg\beta\Rightarrow$, by ($\neg\Rightarrow$). From this, and by using axiom $\Gamma,\beta\Rightarrow\beta$ once again, by ($\to\Rightarrow$) we infer $\Gamma,\beta,\beta\to\neg\beta\Rightarrow$. Finally, $\Gamma,\beta\to\neg\beta\Rightarrow\neg\beta$ follows by ($\Rightarrow\neg$). In particular, $\Delta_{0},\beta\to\neg\beta\Rightarrow\neg\beta$ is provable in LI and so, by (cutcut) with $\Delta_{0}\Rightarrow\beta\to\neg\beta$, it follows that $\Delta_{0}\Rightarrow\neg\beta$ is provable in LI. That is, $\Delta\vdash_{IPL}\neg\beta$, hence $\neg\beta\in\Delta$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\neg\beta$. Then, there exists a $\neg\beta$-saturated set $\Delta^{\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime}$. By IH, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 8.14, $v_{\Delta^{\prime}}(\neg\beta)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime}$.
(2) $\alpha=\beta\to\gamma$ for some $\beta$ and $\gamma$. By Definition 8.14, $\beta\to\gamma\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\beta\to\gamma$. By (DM), $\Delta,\beta\vdash_{IPL}\gamma$ and so, by (DM) again, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\beta\to\gamma$. Then, there exists a $(\beta\to\gamma)$-saturated set $\Delta^{\prime\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime\prime}$. By IH, $v_{\Delta^{\prime\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 8.14, $v_{\Delta^{\prime\prime}}(\beta\to\gamma)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime\prime}$.
@@
Theorem 8.18 (Completeness of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
@@
Suppose that $\Gamma\nvdash_{IPL}\varphi$. Then, there exists a $\varphi$-saturated set $\Delta$ in LI such that $\Gamma\subseteq\Delta$. By Proposition 8.17, $v_{\Delta}$ is a level valuation in $\mathcal{M}_{IPL}$ such that $v_{\Delta}(\alpha)=\mathbf{T}$ for every $\alpha\in\Gamma$, but $v_{\Delta}(\varphi)\neq\mathbf{T}$. This shows that $\Gamma\not\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Remark 8.19.
Observe that the semantical characterization of $\mathbf{IPL}$ w.r.t. level valuation semantics is quite interesting from a theoretical point of view. However, it has little interest from a practical point of view, given that level valuations are infinitary objects, and it is far from obvious how to check validity of a formula in $\mathbf{IPL}$ by using $\mathcal{R}(\mathcal{M}_{IPL})$. The next step is the definition of an (algorithmic) way to construct intuitionistic truth-tables, based precisely on level valuations.
From soundness and completeness of $\mathbf{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$, several important consequences can be obtained. Some of these consequence will be extremely useful to prove the soundness and completeness of the (non-deterministic) truth-tables decision method for $\mathbf{IPL}$ to be defined in the next section.
Corollary 8.19.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
Proof.
It follows from Theorems 8.18, 8.12 and the fact that LI is finitary. ∎
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. Then, $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ if, and only if, $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
Proof.
It follows from Theorems 8.18, 8.12 and the fact that $\bf LI$} satisfies the deduction metatheorem (recall Remark 8.8). ∎
Proposition 8.20.
Let $v\in\mathcal{L}_{IPL}$ and let $\Gamma=\{\alpha\in For(\Omega)\mid v(\alpha)=\mathbf{T}\}$. Then, $\Gamma$ is a closed theory in $\mathbf{IPL}$, i.e.: if $\Gamma\vdash_{IPL}\alpha$ then $\alpha\in\Gamma$.
Proof.
Let $\Gamma=\{\alpha\in For(\Omega)\mid v(\alpha)=\mathbf{T}\}$ and suppose that $\Gamma\vdash_{IPL}\alpha$. By soundness of $\mathbf{IPL}$ w.r.t. level valuation semantics (Theorem 8.12), $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$. Given that $v\in\mathcal{L}_{IPL}$ such that $v\models\Gamma$, it follows that $v\models\alpha$. That is, $v(\alpha)=\mathbf{T}$ and so $\alpha\in\Gamma$. ∎
The following result will be crucial for proving soundness and completeness of intuitionistic truth-tables:
Proposition 8.21.
Let $v\in\mathcal{L}_{IPL}$, and $\alpha\in For(\Omega)$ such that $v(\alpha)=\mathbf{U}$. Then, there exists $w\in\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and $w(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$.
Proof.
$v\in\mathcal{L}_{IPL}$, and $\alpha\in For(\Omega)$ such that $v(\alpha)=\mathbf{U}$. Since, in particular, $v\in Val^{nd}_{+}(\mathcal{M}_{IPL})$ (recall Definitions 8.5 and 8.6), $v(\alpha)=\mathbf{U}$ implies that $\alpha$ cannot be a propositional variable. By definition of the conjunction and disjunction tables in $\mathcal{M}_{IPL}$ (Figure 3), and since $v(\alpha)=\mathbf{U}$, it follows that $\alpha$ cannot be either a conjunction or a disjunction. Let $\Gamma=\{\gamma\in For(\Omega)\mid v(\gamma)=\mathbf{T}\}$. Then, $\alpha\not\in\Gamma$. By the observations above, there are only two cases to consider:
(1) $\alpha=\neg\beta$ for some $\beta$. As noted above, $\alpha\not\in\Gamma$, i.e., $\neg\beta\not\in\Gamma$. Hence $\Gamma\nvdash_{IPL}\neg\beta$, given that $\Gamma$ is closed by Proposition 8.20. If $\beta\in\Gamma$ then $v(\beta)=\mathbf{T}$ and so, by definition of $\mathcal{M}_{IPL}$, $v(\alpha)=v(\neg\beta)=\mathbf{F}$, a contradiction. Then, $\beta\not\in\Gamma$. Suppose that $\Gamma,\beta\vdash_{IPL}\neg\beta$. As in the proof of Proposition 8.17 we infer that $\Gamma\vdash_{IPL}\neg\beta$. But this lead us to a contradiction since, as pointed out before, $\Gamma\nvdash_{IPL}\neg\beta$. From this, $\Gamma,\beta\nvdash_{IPL}\neg\beta$. Thus, there exists a $\neg\beta$-saturated set $\Delta$ in $\mathbf{IPL}$ such that $\Gamma\cup\{\beta\}\subseteq\Delta$. By Proposition 8.17, the function $v_{\Delta}$ is a level valuation such that $v_{\Delta}(\alpha)=\mathbf{F}$ (since $v_{\Delta}(\beta)=\mathbf{T}$) and $v_{\Delta}(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$, as required.
(2) $\alpha=\beta\to\delta$ for some $\beta$ and $\delta$. Since $\Gamma,\delta\vdash_{IPL}\beta\to\delta$ and $\Gamma\nvdash_{IPL}\beta\to\delta$ then $\Gamma\nvdash_{IPL}\delta$. Suppose that $\Gamma,\beta\vdash_{IPL}\beta\to\delta$. By (DM), $\Gamma,\beta\vdash_{IPL}\delta$ and so, by (DM) again, $\Gamma\vdash_{IPL}\beta\to\delta$, a contradiction. From this, $\Gamma,\beta\nvdash_{IPL}\beta\to\delta$. Thus, there exists a $(\beta\to\delta)$-saturated set $\Delta$ in $\mathbf{IPL}$ such that $\Gamma\cup\{\beta\}\subseteq\Delta$. By reasoning as above, we conclude that $\delta\not\in\Delta$. By Proposition 8.17, $v_{\Delta}\in\mathcal{L}_{IPL}$ is such that $v_{\Delta}(\alpha)=\mathbf{F}$ (since $v_{\Delta}(\beta)=\mathbf{T}$ and $v(\delta)\neq\mathbf{T}$) and $v_{\Delta}(\gamma)=\mathbf{T}$ for every formula $\gamma$ such that $v(\gamma)=\mathbf{T}$, as required.
9 A truth-tables procedure for $\mathbf{IPL}$
After obtaining a suitable RNmatrix for $\mathbf{IPL}$, the next step is the definition of a decision procedure from this. As in the case of S4, this procedure is given by means of branching truth-tables with an algorithm for deleting inadequate rows.
9.1 Intuitionistic partial (level) valuations and intuitionistic truth-tables